OpenStudy (anonymous):

Need help with adding 2 fractions. w/4 + w/3 = 28

OpenStudy (mertsj):

Multiply every term by 12. This will clear all fractions.

OpenStudy (anonymous):

by 12? I am confused.

OpenStudy (mertsj):

\[\frac{w}{4}\times 12 + \frac{w}{3}\times12=28\times12\]

OpenStudy (mertsj):

Now cancel the 4 into the 12 and cancel the 3 into the 12

OpenStudy (anonymous):

so 3w + 4w=336 is my new equation?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

ok cool thanks a bunch

OpenStudy (mertsj):

yw. did you get w = 48?

OpenStudy (anonymous):

yes i did. Thanks again.

OpenStudy (mertsj):

yw